# Not all Layers of LLMs are Necessary during Inference

### v2

@@ Abstract
The inference phase of Large Language Models (LLMs) is very expensive. An ideal inference stage of LLMs could utilize fewer computational resources while still maintaining its capabilities (e.g., generalization and in-context learning ability). In this paper, we try to answer the question, “During LLM inference, can we use shallow layers for easy instances; and deep layers for hard ones?” To answer this question, we first indicate that Not all Layers are Necessary during Inference by statistically analyzing the activated layers across tasks. Then, we propose a simple algorithm named AdaInfer to determine the inference termination moment based on the input instance adaptively. More importantly, AdaInfer does not alter LLM parameters and maintains generalizability across tasks. Experiments on well-known LLMs (i.e., Llama2 series and OPT) show that AdaInfer saves an average of 14.8% of computational resources, even up to 50% on sentiment tasks, while maintaining comparable performance. Additionally, this method is orthogonal to other model acceleration techniques, potentially boosting inference efficiency further.
@@
Siqi Fan2, Xin Jiang1, Xiang Li1, Xuying Meng3, Peng Han2, Shuo Shang2*, Aixin Sun4, Yequan Wang1*, Zhongyuan Wang1 1Beijing Academy of Artificial Intelligence, Beijing, China 2University of Electronic Science and Technology of China, Chengdu, China 3Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China 4School of Computer Science and Engineering, Nanyang Technological University, Singapore
@@ 1 Introduction
LLMs have demonstrated impressive performance on various downstream tasks (e.g., text generation, question & answering, and sentiment analysis) using various evaluation protocols such as zero-shot, few-shot, and fine-tuning Todd et al. (2024); Chan et al. (2022); Kossen et al. (2023); Wang et al. (2023, 2022). Notably, In-context learning ability allows LLMs to adapt to tasks using input-output examples without parameter updates Kossen et al. (2023); Todd et al. (2024). However, their inference phases are very expensive Pope et al. (2023); Liu et al. (2023). For example, the inference time complexity for typical large models with Transformer structure is per single inference, where , , and represent the hidden size, sequence length, and layer number, respectively. An ideal inference LLM should utilize fewer computational resources while still maintaining its capabilities in generalization and in-context learning ability Liu et al. (2023). The popular methods for achieving efficient inference in LLMs include model pruning Liu et al. (2018) and sparse models LeCun et al. (1989). However, altering LLM parameters may risk compromising its generalization ability, which is challenging to detect. Meanwhile, different LLM designs pose compatibility challenges with other acceleration methods.
In this paper, we consider dynamically reducing the number of activated neurons as an approach to accelerate LLM inference. Inspired by the human thinking process Salthouse (1996); Deary et al. (2001), where quick answers are often provided for simple questions while more time is spent on thoughtful reasoning for complex ones, e.g., knowledge-related questions. Previous studies Teerapittayanon et al. (2016); Huang et al. (2017) show that “Easy” tasks activate at shallower layers while “hard” ones at deeper layers. Additionally, growth strategy Li et al. (2023) is proposed to lower the training cost of LLMs by adding parameters in stages. It inspires us that reducing the computing parameters during inference may be an effective way besides existing typical accumulation methods. Statistical LLMs results on various tasks (see Section 3.2 for detail) show that reducing parameters is feasible during LLM inference.
@@
Therefore, a natural approach to achieve LLM efficient inference is to decide when to stop the inference process based on the input instance adaptively. For instance, allocating fewer computational resources for processing “simple” samples to enhance operational efficiency. Furthermore, exploring adaptive inference may bridge LLMs with the brain’s information processing Hubel and Wiesel (1962); Murata et al. (2000), aiding in the analysis of activated network modules during sample processing Han et al. (2021) and identifying crucial input components affecting the final prediction.
Specifically, we present AdaInfer, a simple but effective algorithm for instance-aware adaptive inference. The core of AdaInfer lies in data-driven decision-making. Generally, there are two approaches to getting decision-making signals: (1) updating LLM parameters requires training, involves high costs, and might decrease the model’s generalizability Gu et al. (2024), and (2) keeping parameters unchanged, a more desirable and cost-effective approach that preserves the model’s innate ability Yao et al. (2023). In this work, we adopt an early stop strategy, optimizing efficiency without altering the model’s parameters. In particular, we begin by performing statistical analysis on LLM for each block feature (e.g., logits, hidden state, mlp, and attention activation value). Subsequently, we choose logits to construct features and employ classical statistical classifiers (i.e., SVM and CRF) to facilitate the early exit strategy (Section 4).
Experiments on well-known LLMs (i.e., Llama2 series and OPT) show that AdaInfer can save an average of 14.8% computational resources, even up to 50% on sentiment tasks while maintaining comparable performance. More importantly, AdaInfer is orthogonal to other model acceleration techniques, offering the potential for further enhancing inference efficiency (Section 5).
@@ 2 Related Work
A straightforward approach to achieve adaptive inference involves dynamic neural networks Han et al. (2021); Huang et al. (2017); Bolukbasi et al. (2017). Consequently, networks with dynamic structures can be classified into two types: dynamic depth (number of network layers) and width (number of channels, parallel subnetworks etc.).
Dynamic depth.
Dynamic depth involves two methods: Early Exit (EE) and Skip layer. EE first appeared in CNN/DNN networks for visual tasks Bolukbasi et al. (2017); Huang et al. (2017); Teerapittayanon et al. (2016). Subsequently, it was utilized in accelerating the inference of encoder-only architectures in BERT by Li et al. (2020); Liu et al. (2020); Li et al. (2021); Kong et al. (2022). Recently, Schuster et al. (2022); Varshney et al. (2023) discuss confidence-based EE for LM adaptive inference. Our proposed AdaInfer closely aligns with EE concept. Specifically, we apply EE to mainstream decoder-only LLMs, which adhere to the scaling law but suffer from high inference costs due to their large parameter count. Meanwhile, skip-layer dynamically omits the execution of middle layers (or modules) for any input token, facilitated by a gate function Wang et al. (2018) or a binary router Zeng et al. (2023) and layer pruning Kim et al. (2024); Yang et al. (2024); Song et al. (2024). The main difference between our method and theirs is that we achieve instance-wise inference without altering the model parameters, which is crucial for current LLMs. To the best of our knowledge, this is the first attempt to discover that each block’s logits are crucial elements for EE classifiers in LLMs, and we incorporate it as a fundamental design choice in AdaInfer.
Dynamic width.
Dynamic width controls the number of neurons on the network width for efficient inference, such as reducing the number of CNN channels Hua et al. (2019); Hoefler et al. (2021) and establishing multiple parallel structures for “experts” in MoE Fedus et al. (2022); Zhou et al. (2022); Artetxe et al. (2021), dynamically weighting and predicting the output results. Recently, Ma et al. (2023); Addanki et al. (2023); Xia et al. (2023) slimming the network width by pruning attention heads and the output neuron in Query, Key or Value. Other model accelerate methods such as quantization Xiao et al. (2023); Frantar et al. (2022), sparsity Liu et al. (2023) are orthogonal areas and usually excel in different settings.
@@ 3 Efficiency Analysis of LLM Inference
This section aims to prove that Not all Layers are Necessary during Inference by analyzing the number of activated layers across various tasks. We first briefly review LLM’s critical components. Then, we present our statistical observations and insights.
@@ 3.1 Preliminary: LLM Building Blocks
Modern LLMs, rooted in the Transformer architecture Vaswani et al. (2017), are trained with different unsupervised training objectives. For instance, mainstream LLMs (e.g., GPT, Llama series) are pre-trained with a full language modeling objective with a decoder-only structure, computing loss on all tokens. The key components of LLMs can be broken down into the following blocks:
Tokenizer and Embedding Layer.
This block tokenizes input text into numerical vectors, enabling effective processing and analysis of textual data.
Decoder Block.
This block processes numerical vectors through self-attention and feedforward neural networks, enabling the model to focus on (attend to) the most relevant input parts.
Classification Layer.
The LM head layer converts decoder logits into a vocabulary-wide probability distribution to facilitate word prediction.
These blocks facilitate LLMs in efficiently handling NLP downstream tasks, with a primary emphasis on decoder blocks within multi-layer Transformers. For typical large Transformer models, inference complexity is linearly related to the number of decoder layers and is given by per single inference. Consequently, to explore the possibility of skipping intermediate layers in LLMs during inference, we do the following statistics.
@@ 3.2 Not all Layers are Necessary
Earlier Transformer models typically comprise 6 decoder layers, while current open-source models, such as Llama2-13B Touvron et al. (2023), feature 40 decoder layers. However, during inference, each input instance for different tasks passes through every block layer by layer until the last layer, prompting us to question: “Can we allocate fewer computational resources per input instance instead of the same substantial budget?” To investigate this, we conduct a statistical analysis to examine the correlation between accuracy and the activation of layers across various tasks. The statistical results are depicted in Figure 1.
Observation 1:
Not all Layers of LLMs are Necessary during Inference: Early Stopping works. In sentiment analysis using the Llama2-13B (40 layers) model, the average activated layer count per input is , with a variance of . This observation is intuitive. For instance, simpler inputs like “I like Camera A” activate 16 layers, while more complex inputs like “Camera A is better than Camera B in picture quality” activate 24 layers. The latter sentence introduces a comparative sentiment about the “quality” aspect between Camera A and Camera B, which embodies more complex features, suggesting deeper layers for such complex instances.
Observation 2:
Varying Task Difficulties, Different Activation Layers: Stop Simpler Tasks Sooner, Let Complex Ones Go Deeper. Tasks in the LLM activate different layers, with simpler ones usually at shallower layers and more complex ones at deeper layers. This is shown in Figure 1, which demonstrates the performance of a Llama2-7B model across 32 layers in sentiment analysis Socher et al. (2013) and MMLU Hendrycks et al. (2021). For simple tasks like sentiment classification, accuracy matches that of the final layer by the 24th layer. Conversely, for complex tasks like MMLU, accuracy tends to improve with deeper layers.
@@
The observations mentioned above are intuitive. It’s worth noting that similar observations have been made by Teerapittayanon et al. (2016); Huang et al. (2017) for visual tasks in convolutional neural networks. Surprisingly, we have also observed this phenomenon at LLM inference time. By exploiting this phenomenon, we can perform instance-aware adaptive inference for LLMs, dynamically adjusting their structure/parameters for different test samples, thereby achieving superior advantages in inference efficiency and adaptability. Moving forward, we will leverage this observation to implement adaptive inference.
4 AdaInfer
The workflow of AdaInfer and the computational efficiencies gained through this method are depicted in Figures 2(a) and 2(b), respectively. The key of AdaInfer is how to find the early stop signal while keeping the original abilities of LLMs. AdaInfer dynamically computes the stopping signal by evaluating critical features (i.e., “gap” and “top prob”). This process involves two main components: a Feature Selection module and a Classifier module. At each layer, the Feature Selection crafts a feature vector for the current input instance. Subsequently, the Classifier (often SVM or CRF for their effectiveness) assesses the stopping signal’s strength. A strong enough signal triggers an early process termination, allowing for the bypass of subsequent decoder layers.
@@ 4.1 Feature Selection
As we mentioned before, modifying LLM parameters requires training and incurs high costs. More importantly, it may pose a potential risk of compromising the model’s generalization capabilities and detecting these issues can be challenging Gu et al. (2024). Hence, we embrace a more desirable and cost-effective approach that preserves the model’s innate abilities without altering parameters. AdaInfer utilizes specially designed features (e.g., “gap” and “top prob”), leveraging a statistical classifier for evaluation stopping signal.
Problem: The lack of features for decision-making.
LLMs capture coarse-grained features in their initial layers and develop more detailed, fine-grained representations in subsequent, deeper layers, facilitated by repeated application of multi-head attention mechanisms and the use of residual connections. However, there is a lack of universal-level features to demonstrate that shallow-level representation is sufficient for the current task. Furthermore, these features need to be inherently universal to ensure compatibility across various LLMs.
Solution: Logits reflect mutation.
To address this, we conducted a visual analysis of diverse features across the layers within each block of LLMs. Our examination focused specifically on:
Gap: Measures the current block’s prediction confidence for the next token, defined as , where represents the probability distribution generated by the current block.
Top Prob: Indicates , the probability estimation by the current block for the most likely next token.
Cosine Similarity: Calculated to evaluate the similarity between the features of current and previous block, including attention activation value (attn), multi-layer perceptron outputs (mlp), and hidden states.
These analyses are showcased in Figure 3. In this figure, we observe the following trends: (1) For Llama2-13B with 40 layers Touvron et al. (2023) across sentiment and MMLU tasks, the “gap” and “top prob” gradually increase during the inference phase, stabilizing in the deeper layers. (2) The activation of “gap” and “top prob” varies across layers for different tasks. These phenomenons are also evident in the Llama2-7B, OPT-13B Zhang et al. (2022), and GPT-J Wang and Komatsuzaki (2021) (See Appendix C). This demonstrates “gap” and “top prob” can serve as universal features, indicating the stopping signal. Notably, these two values remain consistent across diverse tasks, suggesting a versatile classifier applicable to various tasks. Factor study in subsequent experiments also shows that other features (e.g., Cosine Similarity) exhibit subtle differences across layers.

### v3

@@ Abstract
Due to the large number of parameters, the inference phase of Large Language Models (LLMs) is resource-intensive. However, not all requests posed to LLMs are equally difficult to handle. Through analysis, we show that for some tasks, LLMs can achieve results comparable to the final output at some intermediate layers. That is, not all layers of LLMs are necessary during inference. If we can predict at which layer the inferred results match the final results (produced by evaluating all layers), we could significantly reduce the inference cost. To this end, we propose a simple yet effective algorithm named AdaInfer to adaptively terminate the inference process for an input instance. AdaInfer relies on easily obtainable statistical features and classic classifiers like SVM. Experiments on well-known LLMs like the Llama2 series and OPT, show that AdaInfer can achieve an average of 17.8% pruning ratio, and up to 43% on sentiment tasks, with nearly no performance drop (<1%). Because AdaInfer does not alter LLM parameters, the LLMs incorporated with AdaInfer maintain generalizability across tasks.
@@
Siqi Fan2, Xin Jiang1, Xiang Li1, Xuying Meng3, Peng Han2, Shuo Shang2*, Aixin Sun4, Yequan Wang1*, Zhongyuan Wang1 1Beijing Academy of Artificial Intelligence, Beijing, China 2University of Electronic Science and Technology of China, Chengdu, China 3Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China 4College of Computing and Data Science, Nanyang Technological University, Singapore
@@ 1 Introduction
LLMs have demonstrated impressive performance on various downstream tasks using evaluation protocols such as zero-shot, few-shot, and fine-tuning Todd et al. (2024); Chan et al. (2022); Kossen et al. (2023); Wang et al. (2023, 2022). Example applications include text generation, question answering, and sentiment analysis. Notably, the in-context learning ability allows LLMs to adapt to various different tasks using input-output examples without parameter updates Kossen et al. (2023); Todd et al. (2024). However, the inference phases of LLMs are typically very expensive due to their large number of parameters Pope et al. (2023); Liu et al. (2023). Specifically, the inference time complexity for typical large models with a Transformer structure is per single inference, where , , and represent the number of layers, sequence length, and hidden size, respectively.
Existing solutions to achieve more efficient inference in LLMs include model pruning Ma et al. (2023); Kim et al. (2024) and sparse models LeCun et al. (1989); Liu et al. (2023). Both solutions alter LLM parameters and may risk compromising generalization ability. Additionally, different LLM designs pose compatibility challenges with other acceleration methods. Hence, an ideal efficient LLM inference should use fewer computational resources while maintaining generalization and in-context learning abilities Liu et al. (2023).
If we draw an analogy between LLM inference and the human thinking process Salthouse (1996); Deary et al. (2001), where simple questions can be answered quickly and complex questions require more time for reasoning, we may expect LLMs not to use the same inference power to handle all tasks. Teerapittayanon et al. (2016); Huang et al. (2017) show that "easy" tasks activate at shallower layers while "hard" ones activate at deeper layers. For LLM training, a growth strategy Li et al. (2023) adds parameters in stages to reduce the overall training cost, i.e., not all training instances use the same set of parameters. Hence, we consider that adjusting the parameters during inference based on the difficulty level of a task may be an effective way for efficient inference.
@@
To this end, we conduct a statistical analysis to examine the correlation between the results obtained in intermediate layers and those in the final layers across various tasks. We made two observations: (i) not all layers of LLMs are necessary during inference, i.e., early stopping works, and (ii) simpler tasks require fewer layers, while more complex tasks require more layers of inference. The key to achieving efficient LLM inference then becomes when to stop the inference process adaptively based on the input instance. Interestingly, exploring adaptive inference may bridge LLMs with the brain’s information processing Hubel and Wiesel (1962); Murata et al. (2000), aiding in the analysis of activated network modules during sample processing Han et al. (2021) and identifying crucial input components that affect the final prediction.
In this paper, we present AdaInfer, a simple yet effective algorithm for instance-aware adaptive inference. The core of AdaInfer lies in data-driven decision-making. We begin by performing a statistical analysis on each block feature of LLMs, such as logits, hidden states, mlp, and attention activation values. Consequently, we choose logits to construct features and employ classical statistical classifiers, SVM and CRF, to predict the optimal layer at which to stop the inference. Experiments on well-known LLMs (i.e., Llama2 series and OPT) show that AdaInfer can achieve an average of 17.8% pruning ratio, and up to 43% on sentiment tasks, with nearly no performance drop (<1%). The cost of collecting the small set of statistical features and running AdaInfer is negligible compared to the cost of LLM inference.
AdaInfer is an early stop strategy that optimizes efficiency without altering the model’s parameters. Therefore, AdaInfer does not affect the model’s generalization and in-context learning abilities. Furthermore, being orthogonal to other model acceleration techniques, AdaInfer offers the potential for further enhancing inference efficiency.
@@ 2 Related Work
Existing solutions for achieving adaptive inference involve dynamic neural networks Han et al. (2021); Huang et al. (2017); Bolukbasi et al. (2017). These solutions can be broadly classified into two groups: dynamic depth (number of network layers) and dynamic width (number of channels, parallel subnetworks, etc.).
Dynamic Depth
involves two methods: Early Exit (EE) and Skip Layer. EE first appeared in CNN/DNN networks for visual tasks Bolukbasi et al. (2017); Huang et al. (2017); Teerapittayanon et al. (2016). Subsequently, it was utilized to accelerate the inference of encoder-only architectures in BERT Li et al. (2020); Liu et al. (2020); Li et al. (2021); Kong et al. (2022). Recently, Schuster et al. (2022); Varshney et al. (2023) discussed confidence-based EE for LM adaptive inference. Our proposed AdaInfer closely aligns with the EE concept. We apply EE to mainstream decoder-only LLMs, which adhere to the scaling law but suffer from high inference costs due to their large parameter count.
Skip Layer dynamically omits the execution of middle layers (or modules) for an input token, facilitated by a gate function Wang et al. (2018), a binary router Zeng et al. (2023); Raposo et al. (2024), or layer pruning Kim et al. (2024); Yang et al. (2024); Song et al. (2024); Men et al. (2024); Ma et al. (2023). The main difference between our method and theirs is that we achieve instance-wise inference (i.e., dynamic pruning ratio tailored to specific tasks) without altering the model parameters, which is crucial for current LLMs. To the best of our knowledge, this is the first attempt to discover that each block’s logits are crucial elements for EE classifiers in LLMs, and we incorporate this insight as a fundamental design choice in AdaInfer.
Dynamic Width
controls the number of neurons in the network width for efficient inference. This includes methods such as reducing the number of CNN channels Hua et al. (2019); Hoefler et al. (2021) and establishing multiple parallel structures for "experts" in Mixture of Experts (MoE) Fedus et al. (2022); Zhou et al. (2022); Artetxe et al. (2021), dynamically weighting and predicting the output results. Recently, Ma et al. (2023); Addanki et al. (2023); Xia et al. (2023) have slimmed the network width by pruning attention heads and the output neurons in Query, Key, or Value. Other model acceleration methods, such as quantization Xiao et al. (2023); Xing et al. (2023) and sparsity Liu et al. (2023); Frantar and Alistarh (2023), are orthogonal areas and usually excel in different settings.
@@ 3 Efficiency Analysis of LLM Inference
Before presenting the statistical observations and insights on LLM inference, we first briefly review LLM’s critical components.
@@ 3.1 Preliminary: LLM Building Blocks
Modern LLMs, rooted in the Transformer architecture Vaswani et al. (2017), can be trained with various unsupervised training objectives. In this paper, we focus on mainstream LLMs like GPT and the Llama series. These models are built with a decoder-only structure and are pre-trained with a full language modeling objective, computing loss on all tokens. Their key components can be broken down into the following blocks: Tokenizer and Embedding Layer, Decoder Block, and Classification Layer. The tokenization and embedding layer converts input text into numerical vectors, enabling effective processing and analysis of textual data. The decoder block processes numerical vectors through self-attention and feedforward neural networks, allowing the model to focus on, or attend to, the most relevant parts of the input. Lastly, the classification layer, or the LM head layer, maps decoder logits into a vocabulary-wide probability distribution to facilitate word prediction. These blocks facilitate LLMs in efficiently handling NLP downstream tasks, with a primary emphasis on the decoder block.
During inference, each input instance passes through the decoder block, layer by layer, until the last layer. Hence, the inference complexity is linearly related to the number of decoder layers in the decoder block. The decoder block of earlier models typically comprised 6 layers, whereas current open-source models have many more. For example, Llama2-7B has 32 layers and Llama2-13B features 40 decoder layers Touvron et al. (2023).
@@ 3.2 Not all Layers are Necessary
To explore the possibility of skipping some intermediate layers during inference, we conduct experiments on two tasks: sentiment analysis Socher et al. (2013) and MMLU Hendrycks et al. (2021). We examine the accuracies obtained by stopping inference at each decoding layer and compare them with the final results, i.e., without stopping inference. The experiments were conducted on both Llama2-7B (32 layers) and Llama2-13B (40 layers), and the same observations hold.
Observation 1.
Not all layers of LLMs are necessary during inference: Early Stopping works.
Using the SST-2 dataset Socher et al. (2013), we conduct sentiment classification experiments on the Llama2-13B (40 layers) model. We perform inference at each layer with a batch size of 1 and record the results. On average, an early exit at layer 21 (with a variance of 5.1) achieves comparable accuracy to the final layer output. Interestingly, simpler inputs like ‘I like Camera A’ activate only 18 layers, while more complex inputs like ‘Camera A is better than Camera B in picture quality’ activate about 24 layers. Early stop works on the Llama2-7B model as well.
Observation 2.
Simpler tasks require fewer layers for inference, while complex tasks go deeper.
Figure 1 plots the accuracies by stopping inference at different decoding layers on a Llama2-7B. For the task of sentiment analysis, the accuracy matches that of the final layer by the 24th layer, represented by solid lines in the figure. For MMLU, a complex task, accuracy tends to improve with deeper layers. A similar trend holds across all four tested settings, from 0-shot to 3-shot learning.
@@
Both observations are intuitive and, in fact, not new. Similar findings have been made in visual tasks with convolutional neural networks Teerapittayanon et al. (2016); Huang et al. (2017) and sentence classification with BERT Liu et al. (2020). We extend these observations to decoder-only LLM inferences.
Based on the two observations, we understand that (i) early stopping works, allowing us to reduce inference costs by stopping at certain decoding layers without compromising model accuracy, and (ii) the number of optimal decoding layers for inference is instance-dependent. The number of optimal decoding layers varies across tasks and even across instances of the same task. Recall the two example sentences for sentiment analysis discussed in Observatoin 1. This means that the layer at which inference stops must be dynamically determined (or predicted) for each input instance.
4 AdaInfer: Adaptive Inferences
Modifying LLM parameters may require additional training and pose a potential risk of compromising the model’s generalization capabilities Gu et al. (2024). In designing AdaInfer, we embrace a cost-effective approach that preserves the model’s innate abilities without altering its parameters. The main idea is to capture signals at each decoding layer and make predictions on whether to stop the inference at the current layer.
The workflow of AdaInfer is depicted in Figure 2(a) with three example input instances. At each decoding layer, a Feature Selection component crafts a feature vector for the current input instance. A binary Classifier then predicts whether to stop the inference, i.e., bypass subsequent decoder layers.
@@ 4.1 Feature Selection
LLMs capture coarse-grained features in their initial layers and develop more detailed, fine-grained representations in deeper layers. This process is facilitated by the repeated application of multi-head attention mechanisms and the use of residual connections. However, there is a lack of features to demonstrate at which stage the representation is sufficient for the current task. Furthermore, these features need to be inherently universal to ensure compatibility across various LLMs.
As a part of feature engineering, we conduct a visual analysis of diverse features from each decoding layer (or decoding block illustrated in Figure 2(a)) of LLMs. Our examination focused specifically on:
Gap measures the current block’s prediction confidence for the next token, defined as , where represents the probability distribution generated by the current block.
Top Prob, , is the probability estimation of the most likely next token by the current block.
Cosine Similarities between the current and the previous blocks, calculated on attention activation value (attn), multi-layer perceptron outputs (mlp), and hidden states, respectively.
Again, we use the sentiment and MMLU tasks on the Llama2-13B (40 layers) model for feature analysis, shown in Figure 3. Observe the following trends: (1) across tasks, both “gap” and “top prob” gradually increase alone the inference phase, stabilizing in the deeper layers. (2) The activation of “gap” and “top prob” varies across layers for different tasks. These phenomenons are also evident in the Llama2-7B, OPT-13B Zhang et al. (2022), and GPT-J Wang and Komatsuzaki (2021) (See Appendix C). The feature analysis suggests that “gap” and “top prob” can serve as universal features for the inference-stopping signal. Notably, these two values remain consistent across two diverse tasks, indicating a versatile discriminating power applicable to various tasks. Factor studies in subsequent experiments also show that other features (e.g., cosine similarities) exhibit subtle differences across layers.